OpenStudy (anonymous):

Slope question. Help please :) Write an equation of a line perpendicular to line CD below in slope-intercept form that passes through the point (-1, 6)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

y = -0.5x + 5.5 y = 0.5x - 5.5 y = -2x - 13 y = 2x + 13

OpenStudy (anonymous):

You have a point which is:\[(-1, 6)\] Remember that points are written as: \[(x, y)\] Now it is the plug and play game: \[y = -0.5x + 5.5\] \[x = -1, y = 6\] \[6 = -0.5(-1) + 5.5\] \[6 = 0.5 + 5.5\] \[6 = 6\] \(6=6\) is true, thus your answer is: A) \(y = -0.5x + 5.5\)

OpenStudy (anonymous):

@LanguageEnthusiast Thank You For Taking Your Time To Explain This! It Really Helped Me. :) Do You Think You Could Help Me Out With a Few More? >.<

OpenStudy (anonymous):

I can try to help you if I can, go ahead and post a new question then tag me when you need help. (Make sure you try solving the question yourself first).